OpenStudy (anonymous):

in young town the population reached 4420 people in 1993. in 1998 there was 5600 residents. whats the population in 2000? (algebra)

OpenStudy (anonymous):

if we consider the increase to be linear, then find the equation of the line o=joining (1993,4420) and (1998,5600) , u r required to find the y value (2000,y) which lies on the line :)

OpenStudy (anonymous):

is the answer 6068?

OpenStudy (anonymous):

it is simple use determinent method to this like form an equation from the given points like 1993x+4420y=0 and 1998x+5600y=0 and 2000x+y=0 now write these equation in matrix form like in AX=B form and find | A | from it then you will get correct answer